10.8-x divided by 5/9=1.8
Answer by checkley79(3341)

(8-X)/(5/9)=1.8 INVERT THE 5/9 & MULTIPLY
(8-X)(9/5)=1.8
(72-9X)/5=1.8 CROSS MULTIPLY
72-9X=5*1.8
-9X=9-72
-9X=-63
X=-63/-9
X=7 ANS.
PROOF:
(8-7)/(5/9)=1.8
1*9/5=1.8
9/5=1.8
1.8=1.8